OpenStudy (anonymous):

Calculate the volume of a square pyramid with a base length of 8 ft and a height of 10 ft. Round your answer to the nearest cubic foot, if necessary.

OpenStudy (anonymous):

The formula for the volume of a pyramid is: \(\dfrac{1}{3}A_b\cdot h\), where \(A_b\) is the area of the base of the pyramid. Since this is a square pyramid, the formula for the volume will be \(\dfrac{1}{3}s^2h\). Can you solve it from there?